OpenStudy (anonymous):

If sec(2Pi/3 +x)=2 find x ?

OpenStudy (anonymous):

\[\sec(2\Pi/3+x)=2 \]

OpenStudy (anonymous):

start with \[\cos(x+\frac{2\pi}{3})=\frac{1}{2}\]then first solve for \[x+\frac{2\pi}{3}\] and then solve for \(x\)

OpenStudy (anonymous):

thanks